OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x^4 - 32x^2 - 144

OpenStudy (anonymous):

help!

OpenStudy (mertsj):

Ok. I see that it is -2i. If -2i is a zero then +2i is also a zero.

OpenStudy (anonymous):

@Mertsj yes!

OpenStudy (mertsj):

That means that x+2i and x-2i are factors

OpenStudy (mertsj):

And if those are factors, then (x+2i)(x-2i)= x^2-4i^2=x^2+4 is a factor. So...

OpenStudy (mertsj):